# Training-time Neuron Alignment through Permutation Subspace for Improving Linear Mode Connectivity and Model Fusion

### v1

@@ Abstract
In deep learning, stochastic gradient descent often yields functionally similar yet widely scattered solutions in the weight space even under the same initialization, causing barriers in the Linear Mode Connectivity (LMC) landscape. Overcoming these barriers is crucial for understanding deep learning dynamics and enhancing model-fusion algorithms. Previous studies highlight the role of permutation symmetry in reducing post-training barriers through network permutation. However, these post-hoc methods, demanding extra computations, are less effective for larger, complex models (e.g., ViT, LLM) due to numerous permutation matrices. Thus, in this paper, we study training-time neuron alignment. Our hypothesis suggests that training-time permutation subspace can reduce LMC barriers for free. We find that pruning at initialization supports this. Beyond pruning, we introduce TNA-PFN, a simple yet lossless algorithm using a partial gradient mask during training. TNA-PFN is theoretically and empirically validated for reducing LMC barriers. It excels in wide model fusion applications, especially in federated learning, where two algorithms based on TNA-FPN, named FedPFN and FedPNU, are proposed to show its prospects even under heterogeneous datasets. Moreover, TNA-PFN can enhance the generalization of model soup (Wortsman et al., 2022) for vision transformers and ColD fusion (Don-Yehiya et al., 2023) for pretrained language models.
1 Introduction
Understanding the loss landscape of Deep Neural Networks (DNNs) is the key to understanding the mechanisms and training dynamics behind generalization and optimization (Li et al., 2018b; Fort & Jastrzebski, 2019; Simsek et al., 2021; Vlaar & Frankle, 2022), and it is still an open problem. Empirical findings demonstrate that Stochastic Gradient Descent (SGD) will find many minima that are functionally similar but far away from each other in parameter space (Draxler et al., 2018; Zhang et al., 2021; Entezari et al., 2022). The literature in Linear Mode Connectivity (LMC) suggests that if we linearly interpolate two independently trained networks, which have the same initialization and trainset but have different SGD random seeds (i.e., batch orders), there will be a loss barrier in the landscape (Draxler et al., 2018; Garipov et al., 2018; Ainsworth et al., 2022) (also see Figure 3). The loss barrier reflects that the two SGD solutions fall into two different loss basins that cannot be linearly connected, and it is detrimental to model fusion (Ainsworth et al., 2022; Li et al., 2022).
Figure 1: A simple demonstration of reducing permutation symmetries. There are 6 networks (1 original + 5 counterparts) which are functionally identical but with different permutations. The number of permutation symmetries can be reduced by asymmetrically fixing some weights (in red). Though this demonstration presents a simple static network permutation, we will show in the main paper that this kind of method can realize better neuron alignment under training dynamics.
The reasons behind the barrier in LMC are mainly the over-parameterization and permutation symmetry (also known as permutation invariance) properties of DNNs. Over-parameterization explains why there are abundant minima found by SGD (Zhang et al., 2021; Neyshabur et al., 2017; Safran & Shamir, 2018). Permutation symmetry suggests that the function of the network can remain the same while changing the permutations of neurons, which can result in many functionally same but geometrically different solutions (Ainsworth et al., 2022; Wang et al., 2020a; Tatro et al., 2020). Previous works seek to achieve neuron alignment for improving LMC by utilizing the permutation symmetry property after training. In Entezari et al. (2022), it is conjectured that if taking all permutations into account, all SGD solutions can be mapped into the same loss basin where no barrier in LMC. Git Re-Basin (Ainsworth et al., 2022) further validates this conjecture by proposing three algorithms to find such permutations.
@@ 1 Introduction
However, post-hoc neuron alignment is a hard combinatorial optimization problem. As stated in (Ainsworth et al., 2022), even for a three-layer MLP with 512 widths, the number of permutation symmetries is nearly 10^3498. When comes to larger foundation models, especially vision transformers (Dosovitskiy et al., 2021) and large language models (Radford et al., 2018), finding such an appropriate permutation to align neurons is more challenging and expensive. For the scenarios where alignment and fusion among multiple models are needed, especially federated learning (Wang et al., 2020b), the cost of the post-hoc alignment methods increases as the number of models.
Can we realize training-time neuron alignment?
If training-time alignment is reached, it will relieve the cost and necessity of post-matching. Also, when the model size and the number of models grow, the training-time method can boost model fusion for free without growing computations. In this paper, we hypothesize the key is to break permutation symmetry via permutation asymmetry subspace. During training, if the potential permutation symmetries are reduced, the neurons will be more aligned in a subspace so that the LMC barriers will decrease and model fusion will improve. A simple demonstration of how permutation asymmetry subspace works is in Figure 1.
We find that pruning at initialization supports our hypothesis. To overcome the limitations of pruning, we introduce TNA-PFN, a simple yet lossless algorithm using a partial gradient mask during training, which is then theoretically and empirically validated for reducing LMC barriers. TNA-PFN excels in wide model fusion applications, including in federated learning under heterogeneous datasets, model soup (Wortsman et al., 2022) for vision transformers, and ColD fusion (Don-Yehiya et al., 2023) for pretrained language models. In a nutshell, we have the following contributions.
@@
  • We discover the neuron alignment problem from the perspective of training time, which provides new insights. We hypothesize that learning in permutation subspaces can reach better LMC.
@@
  • Under the hypothesis, we first find pruning at initialization can improve LMC. Then, we propose a simple yet more lossless training-alignment method TNA-PFN, which is validated under both theoretical and empirical analysis.
@@
  • TNA-PFN excels in wide model fusion applications. We extend TNA-PFN in federated learning and devise two algorithms, FedPFN and FedPNU. Also, TNA-PFN can boost model soup and ColD fusion for free, showcasing its prospects for pretrained foundation models.
2 Background and Related Works
In this section, we provide the basic backgrounds and definitions regarding linear mode connectivity and permutation invariance. Additionally, the preliminary of federated learning is in Appendix D.
@@ 2 Background and Related Works
Linear mode connectivity. In this paper, we focus on the linear mode connectivity of two SGD solutions, which have the same initialization but different data orders [^1] We note that there are other forms of LMC, such as the LMC from the initialization and the trained model (Vlaar & Frankle, 2022), and the LMC between two models with different initializations (Entezari et al., 2022). While in this paper, we focus on LMC and model fusion with specific applications in federated learning, we only consider the LMC cases with the same initialization.. We present the definitions of loss barrier and accuracy barrier below.
Definition 2.1.
@@
(Loss barrier (Entezari et al., 2022)) Let $f_{\mathbf{w}}(\cdot)$ be a function represented by a neural network with parameter vector $\mathbf{w}$ that includes all parameters and $\mathcal{L}(\mathbf{w})$ be the any given loss (e.g., train or test error) of $f_{\mathbf{w}}(\cdot)$. Given two independently trained networks $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, let $\mathcal{L}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})$, for $\alpha\in[0,1]$ be the loss of the linearly interpolated network. The loss barrier $B_{loss}(\mathbf{w}_{1},\mathbf{w}_{2})$ along the linear path between $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ is defined as the highest difference between the loss of the interpolated network and linear interpolation of the loss values of the two networks:
@@
The loss barrier of the above definition is not bounded. To better depict and compare the barrier changes, we then propose a definition of the accuracy barrier which is bounded within $[0,1]$.
Definition 2.2.
@@
Permutation invariance. Permutation invariance refers to the property that the positions (i.e., permutations) of neurons of a given network can be changed without changing the network’s function, and it is also known as permutation symmetry (Ainsworth et al., 2022). We take a multi-layer MLP as an example to demonstrate the property.
@@
The connection between LMC and permutation invariance. In Entezari et al. (2022); Ainsworth et al. (2022), it is conjectured that if applying appropriate permutation matrices to the networks, two SGD solutions that have barriers before can be mapped into the same loss basin and linearly connected (with low barriers). In previous literature, post-hoc matching methods are proposed to approximate the right permutations (Ainsworth et al., 2022; Peña et al., 2023).
The comparison between our paper and Frankle et al. (2020). In Frankle et al. (2020), the authors find that better LMC will result in better results of the lottery ticket hypothesis (a kind of pruning). However, their conclusions cannot infer that pruning will result in better LMC, where our contributions lie (see more detailed discussions in Appendix E). It is noteworthy that we are the first to present this finding and it is non-trivial.
We include detailed related works in Appendix E.
3 Permutation Subpace and Linear Mode Connectivity
3.1 Hypothesis and Preliminary Finding
@@
Network pruning improves LMC. One straight-forward method which fits Hypothesis 3.1 is network pruning at initialization. We apply random weight pruning to an initialized model, and different neurons will have different pruning structures (i.e., pruning ratios and pruning positions), so the symmetric structure of the network is broken, and the permutations of neurons are limited. From this pruned initialization, different training runs will learn in the same subspace where pruned weights are always zeros and permutations of neurons are constrained.
It is validated in Figure 2 that pruning can actually improve LMC compared with vanilla training, which supports our hypothesis. But when the pruning ratio $\rho$ is high (i.e., 0.8 and 0.9), pruning will result in an untrainable network with nearly zero generalization. It is more severe when the initialized model is pretrained transformers, from Figure 14, it can be seen that even if the pruning ratio is mild ($\rho=0.2$), random pruning is fatal to ViT’s network generalization. Thus, to overcome pruning’s drawbacks, we propose a simple yet lossless (in terms of generalization) training-time alignment method in subsection 3.2.
3.2 Training-time Neuron Alignment with Partially Fixed Neurons
3.2.1 Method Formulation
Instead of setting the weights as zeros like pruning, we propose to fix some neurons’ weights, which will also break the network symmetry to reduce the permutations while preserving the expressiveness power of networks. An intuitive demonstration is in Figure 1, by fixing some weights of neurons, the number of potential permutations decreases.
Specifically, given an initial network parameterized by a weight vector $\mathbf{w}^{0}\in\mathbb{R}^{d}$. For $\mathbf{w}^{0}$, we randomly generate a mask for each layer according to the mask ratio $\rho$ (refers to the proportion of zeros in the mask $\mathbf{m}^{0}$), and the whole mask is $\mathbf{m}^{0}\in\{0,~{}1\}^{d}$. In $\mathbf{m}^{0}$, 00 for fixed and $1$ for updated, indicating the parameter update status. We individually train $n$ models with different batch orders or datasets. We set each model’s initialization as $\mathbf{w}_{i}\leftarrow\mathbf{w}^{0},i\in[n]$. Each model $\mathbf{w}_{i},i\in[n]$ conducts the following updates in every SGD iteration:

### v2

# Improving Model Fusion by Training-time Neuron Alignment with Fixed Neuron Anchors

@@ Abstract
Model fusion aims to integrate several deep neural network (DNN) models’ knowledge into one by fusing parameters, and it has promising applications, such as improving the generalization of foundation models and parameter averaging in federated learning. However, models under different settings (data, hyperparameter, etc.) have diverse neuron permutations; in other words, from the perspective of loss landscape, they reside in different loss basins, thus hindering model fusion performances. To alleviate this issue, previous studies highlighted the role of permutation invariance and have developed methods to find correct network permutations for neuron alignment after training. Orthogonal to previous attempts, this paper studies training-time neuron alignment, improving model fusion without the need for post-matching. Training-time alignment is cheaper than post-alignment and is applicable in various model fusion scenarios. Starting from fundamental hypotheses and theorems, a simple yet lossless algorithm called TNA-PFN is introduced. TNA-PFN utilizes partially fixed neuron weights as anchors to reduce the potential of training-time permutations, and it is empirically validated in reducing the barriers of linear mode connectivity and multi-model fusion. It is also validated that TNA-PFN can improve the fusion of pretrained models under the setting of model soup (vision transformers) and ColD fusion (pretrained language models). Based on TNA-PFN, two federated learning methods, FedPFN and FedPNU, are proposed, showing the prospects of training-time neuron alignment. FedPFN and FedPNU reach state-of-the-art performances in federated learning under heterogeneous settings and can be compatible with the server-side algorithm.
I Introduction
Deep neural networks (DNNs) have shown great powers in various machine learning tasks; for instance, large language models [1, 2, 3], represented by GPT-4 [4], represent human-level intelligence in question answering, and diffusion models can generate images or videos that cannot be distinguished from reality [5, 6, 7]. Towards more powerful models, model fusion [8, 9] aims to integrate the knowledge and powers of several DNNs into one model by fusing the model parameters, and it has wide and promising applications [10, 11, 12]. One line of research finds that model fusion in pretrained-finetuned paradigm can improve generalization, such as model soup [12] and ColD fusion [11]. Model soup [12], an application of model fusion, effectively enhances generalization performance by fusing the weights of multiple fine-tuned models, which significantly improves prediction accuracy across various tasks without increasing computational demands during inference. ColD fusion [11] finds that model fusion of fine-tuned language models can be recycled to continually improve the pretrained model they are based upon. In addition, federated learning [13, 14, 15] adopts weighted model fusion on the server to generate global models by integrating the knowledge of local data in a privacy-preserving and communication-efficient manner.
Figure 1: Organization overview of the paper.
Despite the prospects of model fusion, there exist barriers when directly fusing model parameters due to the properties of DNNs. Linear mode connectivity (LMC) studies the fundamental aspects of connectivity in loss landscape between different stochastic gradient descent (SGD) solutions [16, 17, 18, 19], serving as basic theories and foundations for model fusion. The word "connectivity" refers to the landscape of linearly fusing two models as $\mathbf{w}=\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2},\text{s.t.}~\alpha\in[0,1]$. It is found that even if two trained SGD solutions have the same initialization and trainset but have different SGD random seeds (i.e., batch orders), there will be a loss barrier in the LMC [18, 19], not to mention in realistic model fusion cases, models are trained on disjoint and heterogeneous/multi-task datasets [12, 20]. Recent studies find that the barrier in LMC is mainly due to permutation symmetry (also known as permutation invariance) properties of DNNs [18]. Because of permutation symmetry, neurons between models are not aligned, and it is suggested that the function of the network can remain the same while changing the permutations of neurons, which can result in many functionally same but geometrically different solutions. From the perspective of the loss landscape, if neurons are not aligned, different solutions reside in distinct loss basins [19]; directly averaging these models’ parameters will result in a high-loss plateau with poor generalization.
@@ I Introduction
Previous works in LMC try to find the right permutations for post-training neuron alignment. In Entezari et al. [18], it is conjectured that if taking all permutations into account, all SGD solutions can be mapped into the same loss basin where no barrier in LMC. Git Re-Basin [19] further validates this conjecture by proposing three algorithms to find such permutations in a more efficient manner. However, post-training matching has limitations, especially for model fusion with large models. First is the computation complexity that post-hoc neuron alignment is a hard combinatorial optimization problem. As stated in Ainsworth et al. [19], even for a three-layer MLP with 512 widths, the number of permutation symmetries is nearly 10^3498. When comes to larger foundation models, especially vision transformers [21] and large language models [22], finding such an appropriate permutation to align neurons is more challenging and expensive. For the scenarios where alignment and fusion among multiple models are needed, especially federated learning [14, 13], the cost of the post-hoc alignment methods increases as the number of models. Additionally, post-hoc matching requires tailored designs for different modules of DNNs, and one method cannot be applied to all architectures. The early methods, such as simulated annealing [18], git re-basin [19], and OTFusion [23], can be applied to MLP and convolutional layers but are not applicable to attention layers and layer normalizations in the transformer blocks. Though new methods are proposed for transformer fusion [24] later on, whether they are applicable to more new architectures (KAN [25], Mamba [26], etc.) still remains questioned.
This paper, instead of post-training neuron alignment, explores whether training-time neuron alignment can be achieved in an efficient and principal way, which can improve model fusion across various scenarios and model architectures. Also, training-time alignment is orthogonal to post-hoc methods, so it can reduce the cost of post-hoc matching or further improve post-matching if jointly used. In addition, studying training-time neuron alignment can help to understand the behind mechanisms of model parameters and training dynamics, serving as foundations for future research of model fusion [8] and model editing [27].
Towards training-time neuron alignment, in this paper, it is hypothesized that the key is to break permutation symmetry via permutation asymmetry subspace. During training, if the potential permutation symmetries are reduced, the neurons will be more aligned in a subspace so that the LMC barriers will decrease and model fusion will improve. It is found that pruning at initialization supports the hypothesis, but pruning will impair individual model performances. Instead of pruning, this paper introduces TNA-PFN, a simple yet lossless algorithm for training-time alignment. TNA-PFN partially fixes some neuron weights as anchors for reducing the potential of permutation symmetries and training in a subspace. Due to the overparameterization property of DNNs, the neurons are so redundant that fixing a proportion of weights may not impair training performances and may preserve the learned features/knowledge. The hypothesis and TNA-PFN are first proposed and verified in LMC with both theoretical and empirical support since LMC is the simplest form and basic foundation of model fusion. Then, TNA-PFN and the methods derived from TNA-PFN are validated under various model fusion scenarios, including model soup [12] for vision transformers, ColD fusion [11] for pretrained language models, and federated learning under heterogeneous datasets [28, 15, 29]. In a nutshell, this paper has the following contributions.
@@
  • We discover the neuron alignment problem from the perspective of training time, which provides new insights. We hypothesize that learning in permutation subspaces can reach better LMC and model fusion.
@@
  • Under the hypothesis, we first find pruning at initialization can improve LMC. Then, we propose a simple yet more lossless training-alignment method—TNA-PFN, which is validated under both theoretical and empirical analysis.
@@
  • TNA-PFN excels in wide model fusion applications. TNA-PFN can boost model soup and ColD fusion for free, showcasing its prospects for pretrained foundation models. Also, we extend TNA-PFN in federated learning and devise two algorithms, FedPFN and FedPNU, both of which are validated effective for improving the global model’s generalization under extensive experiments.
We study training-time neuron alignment under comprehensive scenarios of model fusion, and the main organization of the paper is in Figure 1. First, in section III, training-time neuron alignment is studied in linear mode connectivity, where hypothesis and preliminary findings are shown. TNA-PFN is proposed to reduce the barriers in LMC and is also verified by theoretical insights. Second, in section IV, TNA-PFN is extended into practice—model fusion under the pretrained-finetuned paradigm, where two settings are considered: model soup of vision transformers and ColD fusion of pretrained language models. Third, in section V, two federated learning algorithms are proposed, which are derived from TNA-PFN. It is validated that the two algorithms can reduce the model drifts caused by data heterogeneity and improve the generalization of global models under extensive settings, such as different datasets, data heterogeneity, local epochs, and client numbers.
II Related Works
II-A Model Fusion
Model fusion is an emerging technique in deep learning that unifies the knowledge of several DNNs into one single model in a cost-effective and data-efficient manner [8, 9]. Broadly speaking, model fusion includes model ensemble [8, 9], which only ensembles the predictions of models instead of parameters [30, 31], while in this paper, we focus on a more focused definition of model fusion that studies the fusion of model parameters [23, 24], also known as model merging [10]. Linear mode connectivity studies the fusion of two SGD solutions (i.e., the modes) [16, 17], which provides theoretical foundations of model fusion. Besides, in federated learning, local models are trained on heterogeneous data, and the server generates a global model in each communication round by fusion of local models’ parameters, where the original model fusion algorithm is called FedAvg [13]. Recently, model fusion techniques have been proposed to improve foundation models under a pretrained-finetuned paradigm. The notion of task arithmetic aims to use model fusion as a technique of model editing by adding or removing task-specific model weight vectors [32, 33]. In addition, this paper focuses on two scenarios of model fusion under the pretrained-finetuned paradigm, namely, model soup and ColD fusion. Model soup [12] enhances the generalization performance of foundation models by fusing the weights of multiple fine-tuned models, which significantly improves prediction accuracy across various tasks without increasing computational demands during inference. ColD fusion [11] finds that model fusion of fine-tuned language models can be recycled to continually improve the pretrained model they are based upon.
II-B Linear Mode Connectivity
Linear mode connectivity refers to the phenomenon that there exists a loss (energy) barrier along the linear interpolation path of two networks, in the cases where i) the two networks have the same initialization and are trained on the same dataset but with different random seeds (data shuffles and augmentations) [19]; ii) the two networks are with different initializations but are trained on the same dataset [18]; iii) the two networks are the initial network and the final trained network [34]. Specifically, Adilova et al. [35] examines the linear mode connectivity of different layers. Valaar et al. [34] studies the relationship between generalization and the initial-to-final linear mode connectivity. Frankle et al. [16] connects linear mode connectivity with the lottery ticket hypothesis and finds better connectivity can result in better pruning performances. Zhao et al. [36] bridges mode connectivity and adversarial robustness. Some works try to extend mode connectivity beyond “linear”, e.g., searching for a non-linear low-loss path [17] or studying mode connectivity under spurious attributes [37].
Permutation invariance (a.k.a. permutation symmetry) refers to the property of neural networks that the positions of neurons can be permuted without changing its function [38, 39], and it is believed to be the primary cause of loss barrier in linear mode connectivity [18, 19]. [18] hypothesizes that if taking the permutation invariance into consideration, all solutions can be mapped into the same low-loss basin with connectivity. Further, Ainsworth et al. [19] validates this hypothesis by using “re-basin” which aims to find the appropriate permutation matrices to map the networks into the same basin. Other methods are also utilized to match the neurons for better model fusion, such as optimal transport [23], Bayesian nonparametric technique [40, 41], Hungarian algorithm [42], graph matching [43], and implicit Sinkhorn differentiation [44]. We note that all these methods are for post-matching after training, while we focus on training-time neuron alignment.
II-C Model Fusion in Federated Learning
Federated learning (FL) is a collaborative training paradigm in which several clients collaboratively train a global model without sharing their data in a communication-efficient and privacy-preserving way [13]. In FL, a central server, as the role of coordinator, collects clients’ local models and generates the global model via model fusion. FedAvg [13] is the typical model fusion method in FL, where the fusion weights of local models are set by the sizes of local datasets. In FedLAW [15] and FedDisco [45], weighted fusion is improved by seeking data-heterogeneity-aware fusion weights or setting the norm of weights smaller than 1 for global weight decay regularization.
In FL, clients have heterogeneous data and conduct several local epochs of training before model fusion, causing model drift of local models. To tackle model drift and improve model fusion, some server-side methods have been proposed in FL. FedDF [46] uses ensemble distillation on the server to improve the model fusion of FedAvg. PFNM [40] and FedMA [14] use Bayesian nonparametric methods for neuron alignment of local models before model fusion. While this paper focuses on training-time alignment methods that are orthogonal to the server-side model fusion techniques. Note that the previous work of PAN [47] shares a similar motivation with this paper’s FL methods. PAN uses position-aware encoding on the data input for aligning intermediate activations, which is orthogonal to the neuron weight anchors in this paper. In addition, PAN only focuses on FL, while this paper studies wide applications of model fusion, such as LMC, model soup, and FL.
II-D Works about the Relationship between Pruning and LMC
In Frankle et al. [16], the authors use linear mode connectivity to study the performances of lottery-ticket-hypothesis-based pruning and find that the sparse pruned model with good connectivity will be more likely to reach the full accuracy after pruning. While in this paper, it is found that random pruning (not necessarily lottery tickets) can improve linear mode connectivity. Though the two papers both discuss the relationship between pruning and linear mode connectivity, they have different focuses and contributions: Frankle et al. [16] finds LMC indicates better results of pruning, whereas we find pruning can improve LMC, and the causal logic is different. It is notable that some concurrent works [48, 49] also find that pruning can improve LMC, but according to the dates of release, this paper may be the first to present the finding, and it is non-trivial.
III Hypothesis, Theory, and Empirical Finding of Training-time Neuron Alignment: a Preliminary Study in Linear Mode Connectivity
As LMC provides basic theories and foundations of model fusion, this section gives a preliminary study about training-time neuron alignment in LMC.
III-A Preliminary of Linear Mode Connectivity
@@
This paper considers the linear mode connectivity of two SGD solutions, which have the same initialization but different data orders [^1] It is noted that there are other forms of LMC, such as the LMC from the initialization and the trained model [34], and the LMC between two models with different initializations [18]. In this paper, LMC servers are preliminaries for model fusion with applications such as federated learning, where different models have the same initialization, so the LMC cases with the same initialization are considered.. The definitions of loss barrier and accuracy barrier are listed below.
Definition III.1.
@@
(Loss barrier [18]) Let $f_{\mathbf{w}}(\cdot)$ be a function represented by a neural network with parameter vector $\mathbf{w}$ that includes all parameters and $\mathcal{L}(\mathbf{w})$ be the any given loss (e.g., train or test error) of $f_{\mathbf{w}}(\cdot)$. Given two independently trained networks $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, let $\mathcal{L}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})$, for $\alpha\in[0,1]$ be the loss of the linearly interpolated network. The loss barrier $B_{loss}(\mathbf{w}_{1},\mathbf{w}_{2})$ along the linear path between $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ is defined as the highest difference between the loss of the interpolated network and linear interpolation of the loss values of the two networks:
@@
The loss barrier of the above definition is not bounded. To better depict and compare the barrier changes, a definition of the accuracy barrier, which is bounded within $[0,1]$, is given below.
Definition III.2.
@@
Permutation invariance. Permutation invariance refers to the property that the positions (i.e., permutations) of neurons of a given network can be changed without changing the network’s function, and it is also known as permutation symmetry [19]. We take a multi-layer MLP as an example to demonstrate the property.
@@
III-B Hypothesis and Preliminary Finding
@@
Network pruning improves LMC. One straight-forward method which fits Hypothesis III.3 is network pruning at initialization. We apply random weight pruning to an initialized model, and different neurons will have different pruning structures (i.e., pruning ratios and pruning positions), so the symmetric structure of the network is broken, and the permutations of neurons are limited. From this pruned initialization, different training runs will learn in the same subspace where pruned weights are always zeros and permutations of neurons are constrained.
It is validated in Figure 2 that pruning can actually improve LMC compared with vanilla training, which supports our hypothesis. But when the pruning ratio $\rho$ (the proportion of eliminated neuron weights) is high (i.e., 0.8 and 0.9), pruning will result in an untrainable network with nearly zero generalization. It is more severe when the initialized model is pretrained transformers, from Figure 7, it can be seen that even if the pruning ratio is mild ($\rho=0.2$), random pruning is fatal to ViT’s network generalization. Thus, to overcome pruning’s drawbacks, we propose a simple yet lossless (in terms of generalization) training-time alignment method in subsection III-C.
III-C Training-time Neuron Alignment with Partially Fixed Neurons as Anchors
III-C1 Method Formulation
Pruning sets the fixed weights as zeros, and multiplying these zero weights will let intermediate features be zeros as $\mathbf{h}_{l}=\mathbf{W}_{l-1}\mathbf{h}_{l-1}$, missing important data features and losing expressiveness of DNNs if pruning is severe. Instead of setting the weights as zeros like pruning, we propose to fix some neurons’ weights, which will also break the network symmetry to reduce the permutations while preserving the expressiveness power of networks. An intuitive demonstration is in Figure 3, by fixing some weights of neurons, the number of potential permutations decreases. The fixed neuron weights can serve as anchors for reducing the potential number of network permutations. The proposed method is called Training-time Neuron Alignment with Partially Fixed Neurons as Anchors, dubbed as TNA-PFN. Due to the overparameterization and redundancy properties of DNNs, partially fixing some weights as TNA-PFN will not impede the optimization, instead, serving as regularization for better generalization.
We give the detailed implementation of TNA-PFN via unified gradient mask. To keep the weights frozen, we use fixed and unified gradient masks to keep certain neuron weights from updating. Specifically, given an initial network parameterized by a weight vector $\mathbf{w}^{0}\in\mathbb{R}^{d}$. For $\mathbf{w}^{0}$, we randomly generate a mask for each layer according to the mask ratio $\rho$ (refers to the proportion of zeros in the mask $\mathbf{m}^{0}$, i.e., the proportion of fixed weights), and the whole mask is $\mathbf{m}^{0}\in\{0,~1\}^{d}$. In $\mathbf{m}^{0}$, 0 for fixed and 11 for updated, indicating the parameter update status. We individually train $n$ models with different batch orders or datasets. We set each model’s initialization as $\mathbf{w}_{i}\leftarrow\mathbf{w}^{0},i\in[n]$. Each model $\mathbf{w}_{i},i\in[n]$ conducts the following updates in every SGD iteration using the identical mask $\mathbf{m}^{0}$ for neuron alignment: